OpenStudy (anonymous):

A 4,000 mL solution of AgNO3 contains 17.00 g of solute in water. Calculate the molarity (molar concentration) of the solution.

OpenStudy (taramgrant0543664):

c=(m/v)x(1/MW) Where c is the molar concentration or molarity, your unknown m is the mass of solute in grams, so the 17.00g v is the volume in litres, so you would have to convert it to L from mL MW is the molecular weight so you have to find the molar mass of AgNO3

OpenStudy (abb0t):

molarity is defined as moles per litre. Meaning, \(\large \sf \color{red}{\frac{mol}{L}}\) Most of what you're doing in general chemistry is following units. If you know the units, you can get to where you need to get.

OpenStudy (abb0t):

Keeping in mind that you need to eliminate grams of solute and convert to moles using molar mass (which is in units of \(\sf \color{red}{\frac{g}{mol}}\). Refer to the periodic table here for this, if you don't have one, here's a nice one you can use: http://www.ptable.com/